Assume that a normal distribution of data has a mean of 14 and a standard deviation of 2. Use the empirical rule to find the percentage of values that lie below 18.
Answer #1

Solution :

Given that,

= 14

= 2

Using Empirical rule,

P( - 2< X < + 2) = 95%

P(14 - 2 * 2 < X < 14 + 2 * 2) = 95%

P(10 < X < 18) = 95%

P(X < 18) = 1 - 0.025 = 0.975

Answer = 97.5%